The sum of the digits of a 2 digit no. is 9, when the digits are reversed the new no. is 27 more than the original no. Find the no.
Answer by venugopalramana(3286)

The sum of the digits of a 2 digit no. is 9, when the digits are reversed the new no. is 27 more than the original no. Find the no.
LET XY BE THE NUMBER...SUM OF DIGITS =X+Y=9.....................I
ITS VALUE =10X+Y
WHEN DIGITS ARE REVERSED WE GET YX...ITS VALUE = 10Y+X
DIFFERENCE BETWEEN THE 2 NUMBERS =27=10Y+X-10X-Y=9Y-9X=9(Y-X)
Y-X=3.............II
EQN.I+EQN.II....
2Y=9+3=12
Y=12/2=6
X=9-Y=9-6=3
THE NUMBER IS 36